OpenStudy (anonymous):

if f(x)={x^2-1, x<3 2ax, x>=3} what is the value of makes every x become continuous

OpenStudy (anonymous):

what value of a is continuous at every x

OpenStudy (bahrom7893):

a function is continuous if the left hand limit of 3 is equal to the right hand limit

OpenStudy (bahrom7893):

so lim x->3 - would be: 3^2-1=9-1=8 lim x->3 + 2*a*3

OpenStudy (bahrom7893):

by the way lim x->3+ (2*a*3), by that i meant lim as x goes to three from the right side

OpenStudy (bahrom7893):

anyway both limits have to be equal so: 2*a*3 = 8 6a = 8 a = 8/6 = 4/3